# Teuken-7B-Base & Teuken-7B-Instruct: Towards European LLMs

### v2

@@ A.1 Data
In this section, we outline the Common Crawl data utilized in our LLM training, detailing the cutoff dates and the data collection period. The dumps span multiple years, with varying distributions of weeks per year (cf. Table 5), which is critical for understanding the temporal coverage of the training data.
@@ A.2 Training Ablation Experiments
We present here the training ablations and our methodology in more detail. Table 6 summarizes the main results, while Figures 6 and 7 display training loss curves of the various ablations against the baseline (“original” in the plots). The loss curves are throughput-normalized where appropriate.
Our baseline for the ablation experiments was the best-performing 2.6B parameter model from (Ali et al., 2024). To make experiments comparable, the training ablations used the same training setup. For evaluation, we focused on 3 tasks: (1) ARC Easy Clark et al. (2018), (2) HellaSwag Zellers et al. (2019), and (3) LAMBADA Paperno et al. (2016). Because the models often achieved better-than-random performance in these evaluation tasks and because they test different abilities of the model, they were deemed a good-enough proxy measure of general downstream improvements.
@@
    SwiGLU: replace the first layer of the MLP part of the Transformer with a T5-style (i.e., without biases) Raffel et al. (2020); Google (2020) Swish-activated Ramachandran et al. (2017) gated linear unit layer Dauphin et al. (2017); Shazeer (2020).
@@
    Untied in/out embedding: learn separate weights for the input embedding and output “unembedding” layers Google (2020).
@@
    No Linear biases: remove all bias terms in Linear layers Google (2020). The ablation ran until 33 000 steps and was not evaluated due to being deemed too far from completion.
@@
    No GPT-like weight init: whether to scale weight initialization in layers that a residual path leads into as a function of depth Radford et al. (2019). The ablation ran until 39 000 steps and was not evaluated due to being deemed too far from completion.
@@
    Head scaling: multiply each Attention head’s output by a learned scalar factor Shleifer et al. (2021).
@@
    No dropout: disable dropout Srivastava et al. (2014) in all layers Google (2020). The latest checkpoint available for evaluation was at 21 000 steps. While it showed a strongly monotonic improvement over the baseline, it is especially hard to interpret improvement in training loss in the dropout vs. no dropout setting. We evaluated this change even though it was far from finishing training because its improvements were so drastic.
@@
    RMSNorm: replace LayerNorm normalization layers Lei Ba et al. (2016) with root mean square layer normalization layers Zhang and Sennrich (2019).
@@
    NoPE: no position embedding Kazemnejad et al. (2024); completely remove position embeddings.
@@
    ALiBi: replace the Rotary position embedding Su et al. (2024) with the Attention with linear biases position embedding Press et al. (2021). Because we were unable to use an optimized ALiBi kernel for this ablation, throughput-normalization was not considered out of fairness.
@@
    GQA (2 groups): replace multi-head Attention with grouped-query Attention Ainslie et al. (2023) with 2 groups (i.e., 2 key/value heads).
@@
    Adan (4× base LR): replace the AdamW optimizer with the Adan optimizer Xie et al. (2024), using the baseline’s learning rate multiplied by 4. The increased learning rate was chosen based on previous small-scale experiments.
@@
Notably, we decided to use neither Adan nor 4× the learning rate despite these changes yielding the best results. Instead, to minimize risks, we opted to use the same learning rate as Llama-2 Touvron et al. (2023b). Due to our inexperience at training at this scale, we were worried about possible convergence problems if we had chosen these important parameters incorrectly.
@@
(a) “RMSNorm” ablation (time-normalized)
(b) “NoPE” ablation (time-normalized)
(c) “ALiBi” ablation (not time-normalized out of fairness)
(d) “GQA (2 groups)” ablation (time-normalized)
(e) “Adan” ablation (time-normalized)
(f) “2×/4× base LR” ablation
@@ A.3 Impact of Educational Content
Figure 8: Downstream performance of the base model from 0.99T to 4T tokens for English. The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens
Figure 9: Downstream performance of the base model from 0.99T to 4T tokens for French. The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens
Figure 10: Downstream performance of the base model from 0.99T to 4T tokens for German. The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens
Figure 11: Downstream performance of the base model from 0.99T to 4T tokens for Finish The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens
Figure 12: Downstream performance of the base model from 0.99T to 4T tokens for Estonian The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens
A.4 Continued Pre-Training
Because we decided to further train our base model from 3T tokens to 4T tokens, we needed to define a new learning rate schedule. We ablated three different learning rate schedules Full-Rewarmup, Partial-Rewarmup, and Continue Decay. In the Full-Rewarmup setting, we defined a cosine learning rate schedule with the same minimal and maximal learning as for the initial pre-training phase. In the Partial-Rewarmup, we set the maximum learning rate to a quarter of the maximum learning rate of the initial pre-training phase. Finally, in the Continue-Decay setting, we decreased the learning rate to 8.3% of the maximum learning rate within the following 1T tokens, providing the lowest training loss (Figure 13). Therefore, we decided to employ this setting to train our model further.
A.5 Evaluation
The tables 9-12 present the results of our base model compared to related base models.
Model Average EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
Aya-23-8B .563 .593 .661 .484 .512
Salamandra-7B-Instruct .557 .620 .664 .476 .468
Bloomz-7B1 .394 .373 .419 .456 .327
Ours (Instruct) .560 .610 .653 .541 .437
Table 8: Instruct model results on multilingual benchmarks across the 10 common languages (Czech, Dutch, English, French, Greek, Italian, Polish, Portuguese, Romanian, and Spanish). The tables include mean accuracy across languages.
@@ A.5 Evaluation
Meta-Llama-3.1-8B .536 .534 .566 .500 .543
Salamandra-7B .518 .580 .627 .455 .412
Mistral-7B-v0.3 .486 .482 .504 .476 .483
Occiglot-7B-eu5 .434 .423 .457 .454 .400
Pharia-1-LLM-7B .365 .321 .360 .457 .322
Bloom-7B1 .333 .284 .316 .477 .254
Ours (Base) .491 .540 .603 .474 .345
@@
Mistral-7B-Instruct-v0.3 .563 .591 .572 .496 .593
Meta-Llama-3.1-8B-Instruct .547 .554 .544 .546 .546
Salamandra-7B-Instruct .545 .595 .642 .449 .493
Mistral-7B-v0.3 .538 .563 .568 .488 .531
Meta-Llama-3.1-8B .534 .540 .559 .516 .520
Salamandra-7B .519 .586 .638 .389 .463
Occiglot-7B-eu5-Instruct .464 .470 .480 .369 .538
Occiglot-7B-eu5 .448 .464 .474 .344 .511
Aya-23-8B .422 .386 .421 .403 .477
Pharia-1-LLM-7B-ctr .335 .264 .305 .270 .500
Pharia-1-LLM-7B-ctr-aligned .333 .253 .310 .263 .507
Bloom-7B1 .332 .259 .295 .259 .517
Bloomz-7B1 .326 .260 .292 .251 .500
Ours (Instruct) .527 .565 .625 .364 .555
Ours (Base) .494 .535 .608 .323 .509
Table 13: ask accuracies for the Bulgarian language.
@@
Ours (Instruct) .540 .588 .618 .423 .530
Ours (Base) .494 .557 .601 .353 .464
@@
Ours (Instruct) .548 .598 .650 .447 .496
Ours (Base) .489 .553 .636 .349 .418
@@
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .591 .610 .621 .615 .520
Mistral-7B-Instruct-v0.3 .584 .620 .609 .540 .568
Meta-Llama-3.1-8B .577 .589 .632 .590 .495
Occiglot-7B-eu5-Instruct .575 .625 .691 .514 .469
Pharia-1-LLM-7B-ctr-aligned .567 .618 .669 .482 .498
Aya-23-8B .564 .589 .645 .519 .501
Occiglot-7B-eu5 .561 .615 .679 .512 .439
Mistral-7B-v0.3 .552 .595 .598 .551 .465
Pharia-1-LLM-7B-ctr .546 .616 .659 .450 .457
Salamandra-7B-Instruct .543 .614 .644 .466 .451
Salamandra-7B .531 .601 .645 .432 .445
Bloomz-7B1 .340 .299 .307 .309 .442
Bloom-7B1 .323 .294 .310 .262 .427
Ours (Instruct) .564 .621 .644 .451 .542
Ours (Base) .507 .577 .634 .367 .452
Table 16: Task accuracies for the German language.
@@
Ours (Instruct) .522 .573 .621 .356 .539
Ours (Base) .486 .540 .611 .315 .479
@@
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Mistral-7B-Instruct-v0.3 .706 .753 .846 .627 .597
Meta-Llama-3.1-8B-Instruct .692 .736 .802 .690 .541
Meta-Llama-3.1-8B .662 .715 .819 .661 .452
Mistral-7B-v0.3 .654 .727 .829 .635 .426
Occiglot-7B-eu5-Instruct .624 .698 .797 .551 .449
Pharia-1-LLM-7B-ctr-aligned .623 .712 .775 .519 .486
Aya-23-8B .617 .670 .779 .565 .454
Salamandra-7B-Instruct .609 .711 .772 .508 .446
Pharia-1-LLM-7B-ctr .607 .702 .770 .503 .454
Occiglot-7B-eu5 .605 .681 .790 .543 .404
Salamandra-7B .586 .695 .771 .453 .423
Bloomz-7B1 .496 .550 .610 .372 .452
Bloom-7B1 .449 .546 .606 .257 .389
Ours (Instruct) .618 .702 .759 .497 .514
Ours (Base) .563 .674 .749 .415 .413
Table 18: Task accuracies for the English language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .618 .641 .673 .627 .530
Mistral-7B-Instruct-v0.3 .602 .655 .654 .548 .552
Meta-Llama-3.1-8B .597 .632 .679 .606 .470
Occiglot-7B-eu5-Instruct .591 .656 .729 .520 .458
Pharia-1-LLM-7B-ctr-aligned .578 .648 .693 .480 .492
Occiglot-7B-eu5 .576 .636 .712 .516 .438
Mistral-7B-v0.3 .574 .628 .651 .561 .456
Salamandra-7B-Instruct .566 .643 .687 .483 .450
Aya-23-8B .566 .608 .677 .523 .456
Pharia-1-LLM-7B-ctr .552 .636 .681 .453 .437
Salamandra-7B .549 .634 .686 .434 .443
Bloomz-7B1 .478 .517 .582 .379 .434
Bloom-7B1 .437 .509 .573 .260 .407
Ours (Instruct) .565 .628 .667 .448 .520
Ours (Base) .513 .591 .665 .354 .442
Table 19: Task accuracies for the Spanish language.
@@
Ours (Instruct) .499 .507 .544 .405 .540
Ours (Base) .453 .484 .536 .342 .452
@@
Bloomz-7B1 .307 .267 .296 .269 .395
Ours (Instruct) .511 .529 .590 .414 .511
Ours (Base) .466 .511 .587 .343 .424
@@
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .612 .638 .667 .624 .521
Mistral-7B-Instruct-v0.3 .606 .647 .660 .549 .567
Meta-Llama-3.1-8B .593 .620 .675 .598 .478
Occiglot-7B-eu5-Instruct .581 .656 .722 .511 .435
Pharia-1-LLM-7B-ctr-aligned .575 .643 .698 .477 .482
Mistral-7B-v0.3 .573 .627 .652 .562 .451
Aya-23-8B .564 .610 .687 .519 .441
Occiglot-7B-eu5 .563 .643 .713 .504 .393
Salamandra-7B-Instruct .555 .635 .681 .461 .444
Pharia-1-LLM-7B-ctr .551 .636 .684 .458 .428
Salamandra-7B .532 .624 .683 .430 .391
Bloomz-7B1 .461 .500 .579 .367 .396
Bloom-7B1 .431 .501 .573 .262 .387
Ours (Instruct) .576 .623 .676 .448 .558
Ours (Base) .519 .596 .665 .357 .457
Table 22: Task accuracies for the French language.
@@
Ours (Instruct) .518 .538 .580 .402 .554
Ours (Base) .470 .521 .568 .299 .489
@@
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .616 .635 .649 .614 .565
Mistral-7B-Instruct-v0.3 .590 .625 .627 .545 .565
Occiglot-7B-eu5-Instruct .589 .647 .708 .519 .483
Meta-Llama-3.1-8B .588 .624 .655 .592 .483
Pharia-1-LLM-7B-ctr-aligned .575 .627 .679 .480 .515
Occiglot-7B-eu5 .575 .629 .702 .517 .451
Aya-23-8B .565 .596 .663 .513 .488
Mistral-7B-v0.3 .561 .601 .618 .556 .471
Salamandra-7B-Instruct .558 .626 .658 .472 .478
Pharia-1-LLM-7B-ctr .554 .616 .670 .457 .473
Salamandra-7B .536 .610 .656 .432 .448
Bloomz-7B1 .368 .339 .380 .326 .428
Bloom-7B1 .363 .345 .390 .263 .453
Ours (Instruct) .566 .604 .649 .448 .562
Ours (Base) .506 .573 .640 .356 .457
Table 24: Task accuracies for the Italian language.
@@
Ours (Instruct) .509 .518 .558 .392 .566
Ours (Base) .471 .506 .557 .317 .502
@@
Ours (Instruct) .512 .507 .557 .408 .577
Ours (Base) .477 .484 .548 .349 .527
@@
Ours (Instruct) .560 .596 .642 .443 .558
Ours (Base) .515 .561 .635 .372 .490
@@
Ours (Instruct) .546 .580 .622 .425 .558
Ours (Base) .497 .546 .603 .350 .490
@@
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .608 .625 .651 .622 .535
Mistral-7B-Instruct-v0.3 .590 .625 .625 .551 .558
Meta-Llama-3.1-8B .587 .624 .662 .596 .466
Aya-23-8B .570 .608 .674 .517 .479
Pharia-1-LLM-7B-ctr-aligned .563 .611 .661 .471 .508
Mistral-7B-v0.3 .562 .614 .624 .561 .449
Salamandra-7B-Instruct .557 .631 .667 .473 .459
Pharia-1-LLM-7B-ctr .542 .603 .649 .448 .468
Occiglot-7B-eu5-Instruct .541 .593 .623 .490 .457
Salamandra-7B .535 .619 .669 .430 .422
Occiglot-7B-eu5 .526 .578 .615 .491 .420
Bloomz-7B1 .455 .490 .555 .373 .403
Bloom-7B1 .433 .490 .553 .255 .434
Ours (Instruct) .554 .614 .658 .450 .496
Ours (Base) .516 .583 .652 .375 .452
Table 29: Task accuracies for the Portuguese language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .583 .589 .589 .596 .558
Meta-Llama-3.1-8B .568 .577 .597 .580 .518
Mistral-7B-Instruct-v0.3 .563 .584 .550 .519 .597
Aya-23-8B .559 .578 .638 .505 .516
Mistral-7B-v0.3 .546 .558 .550 .537 .540
Salamandra-7B-Instruct .546 .598 .629 .465 .492
Salamandra-7B .533 .595 .634 .418 .485
Occiglot-7B-eu5-Instruct .490 .516 .499 .433 .510
Occiglot-7B-eu5 .474 .494 .488 .438 .477
Pharia-1-LLM-7B-ctr-aligned .377 .322 .336 .352 .497
Pharia-1-LLM-7B-ctr .372 .314 .333 .350 .490
Bloomz-7B1 .346 .287 .302 .299 .497
Bloom-7B1 .341 .287 .302 .258 .518
Ours (Instruct) .553 .589 .615 .434 .572
Ours (Base) .508 .558 .605 .355 .513
Table 30: Task accuracies for the Romanian language.
@@
Ours (Instruct) .525 .552 .598 .423 .527
Ours (Base) .481 .524 .590 .354 .454
@@
Ours (Instruct) .527 .555 .577 .418 .558
Ours (Base) .477 .512 .575 .345 .474

### v3

@@ A.1 Data
In this section, we outline the Common Crawl data used, detailing the cutoff dates and the data collection period. The dumps span multiple years, with varying distributions of weeks per year (cf. Table˜4), which is critical for understanding the temporal coverage of the training data.
@@ A.2 Training Ablation Experiments
We present here the training ablations and our methodology in more detail. Table 5 summarizes the main results, while Figures 7 and 14 display training loss curves of the various ablations against the baseline (“original” in the plots). The loss curves are throughput-normalized where appropriate.
We conducted the experiments on 2.6B scale, and focused on 3 downstream-tasks: (1) ARC Easy [40], (2) HellaSwag [44], and (3) LAMBADA [18]. Because the models often achieved better-than-random performance in these evaluation tasks and because they test different abilities of the model, they were deemed a good-enough proxy measure of general downstream improvements.
@@
    SwiGLU: replace the first layer of the MLP part of the Transformer with a T5-style (i.e., without biases) [13, 61] Swish-activated [42] gated linear unit layer [58, 66].
@@
    Untied in/out embedding: learn separate weights for the input embedding and output “unembedding” layers [61].
@@
    No Linear biases: remove all bias terms in Linear layers [61]. The ablation ran until 33 000 steps and was not evaluated due to being deemed too far from completion.
@@
    No GPT-like weight init: whether to scale weight initialization in layers that a residual path leads into as a function of depth [65]. The ablation ran until 39 000 steps and was not evaluated due to being deemed too far from completion.
@@
    Head scaling: multiply each Attention head’s output by a learned scalar factor [49].
@@
    No dropout: disable dropout [38] in all layers [61]. The latest checkpoint available for evaluation was at 21 000 steps. While it showed a strongly monotonic improvement over the baseline, it is especially hard to interpret improvement in training loss in the dropout vs. no dropout setting. We evaluated this change even though it was far from finishing training because its improvements were so drastic.
@@
    RMSNorm: replace LayerNorm normalization layers [30] with root mean square layer normalization layers [12].
@@
    NoPE: no position embedding [5]; completely remove position embeddings.
@@
    ALiBi: replace the Rotary position embedding [27] with the Attention with linear biases position embedding [39]. Because we were unable to use an optimized ALiBi kernel for this ablation, throughput-normalization was not considered out of fairness.
@@
    GQA (2 groups): replace multi-head Attention with grouped-query Attention [23] with 2 groups (i.e., 2 key/value heads).
@@
    Adan (4× base LR): replace the AdamW optimizer with the Adan optimizer [57], using the baseline’s learning rate multiplied by 4. The increased learning rate was chosen based on previous small-scale experiments.
@@
Notably, we decided to use neither Adan nor 4× the learning rate despite these changes yielding the best results. Instead, to minimize risks, we opted to use the same learning rate as Llama-2 [22]. Due to our inexperience at training at this scale, we were worried about possible convergence problems if we had chosen these important parameters incorrectly.
@@
Figure 8: “RMSNorm” ablation (time-normalized)
Figure 9: “NoPE” ablation (time-normalized)
Figure 10: “ALiBi” ablation (not time-normalized out of fairness)
Figure 11: “GQA (2 groups)” ablation (time-normalized)
Figure 12: “Adan” ablation (time-normalized)
Figure 13: “2×/4× base LR” ablation
@@ A.3 Impact of Educational Content
Figures 15-19 demonstrate the development of the model downstream performance between 0.99T and 6T tokens for English, French, German, Finnish, and Estonian. The grey area in the figures highlights the ablation in which we compare the performance of the EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens.
Figure 15: Downstream performance of the base model from 0.99T to 6T tokens for English. The grey area highlights the ablation comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
Figure 16: Downstream performance of the base model from 0.99T to 6T tokens for French. The grey area highlights the ablation comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
Figure 17: Downstream performance of the base model from 0.99T to 6T tokens for German. The grey area highlights the ablation comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
Figure 18: Downstream performance of the base model from 0.99T to 6T tokens for Finish The grey area highlights the ablation comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
Figure 19: Downstream performance of the base model from 0.99T to 6T tokens for Estonian The grey area highlights the ablation study comparing the performance of EU24 data to the FineWeb-EDU dataset between 2.85T and 3T tokens. After 4T tokens, we replaced the English FineWeb-EDU and continued training until 6T tokens with DCLM-Baseline.
A.4 Evaluation
The Tables 7-10 present the results of our base model compared to related base models.
@@ A.4 Evaluation
Meta-Llama-3.1-8B .527 .522 .549 .503 .535
Salamandra-7B .514 .572 .617 .457 .410
Mistral-7B-v0.3 .475 .466 .482 .479 .473
Occiglot-7B-eu5 .422 .406 .435 .458 .391
Pharia-1-LLM-7B .349 .296 .332 .457 .310
Bloom-7B1 .325 .267 .296 .483 .254
Teuken-7B-Base (Ours) .511 .539 .597 .466 .443
@@
Meta-Llama-3.1-8B-Instruct .591 .610 .621 .615 .520
Mistral-7B-Instruct-v0.3 .584 .620 .609 .540 .568
Meta-Llama-3.1-8B .577 .589 .632 .590 .495
Occiglot-7B-eu5-Instruct .575 .625 .691 .514 .469
Pharia-1-LLM-7B-ctr-aligned .567 .618 .669 .482 .498
Aya-23-8B .564 .589 .645 .519 .501
Occiglot-7B-eu5 .561 .615 .679 .512 .439
Mistral-7B-v0.3 .552 .595 .598 .551 .465
Pharia-1-LLM-7B-ctr .546 .616 .659 .450 .457
Salamandra-7B-Instruct .543 .614 .644 .466 .451
Salamandra-7B .531 .601 .645 .432 .445
Bloomz-7B1 .340 .299 .307 .309 .442
Bloom-7B1 .323 .294 .310 .262 .427
Teuken-7B-Instruct (Ours) .588 .607 .689 .477 .577
Teuken-7B-Base (Ours) .530 .576 .636 .470 .436
Table 11: Task accuracies for the German language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Mistral-7B-Instruct-v0.3 .706 .753 .846 .627 .597
Meta-Llama-3.1-8B-Instruct .692 .736 .802 .690 .541
Meta-Llama-3.1-8B .662 .715 .819 .661 .452
Mistral-7B-v0.3 .654 .727 .829 .635 .426
Occiglot-7B-eu5-Instruct .624 .698 .797 .551 .449
Occiglot-7B-eu5 .605 .681 .790 .543 .404
Aya-23-8B .617 .670 .779 .565 .454
Pharia-1-LLM-7B-ctr-aligned .623 .712 .775 .519 .486
Pharia-1-LLM-7B-ctr .607 .702 .770 .503 .454
Salamandra-7B-Instruct .609 .711 .772 .508 .446
Salamandra-7B .586 .695 .771 .453 .423
Bloomz-7B1 .496 .550 .610 .372 .452
Bloom-7B1 .449 .546 .606 .257 .389
Teuken-7B-Instruct (Ours) .635 .685 .820 .516 .517
Teuken-7B-Base (Ours) .581 .667 .771 .512 .372
Table 12: Task accuracies for the English language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .612 .638 .667 .624 .521
Mistral-7B-Instruct-v0.3 .606 .647 .660 .549 .567
Meta-Llama-3.1-8B .593 .620 .675 .598 .478
Occiglot-7B-eu5-Instruct .581 .656 .722 .511 .435
Pharia-1-LLM-7B-ctr-aligned .575 .643 .698 .477 .482
Mistral-7B-v0.3 .573 .627 .652 .562 .451
Aya-23-8B .564 .610 .687 .519 .441
Occiglot-7B-eu5 .563 .643 .713 .504 .393
Salamandra-7B-Instruct .555 .635 .681 .461 .444
Pharia-1-LLM-7B-ctr .551 .636 .684 .458 .428
Salamandra-7B .532 .624 .683 .430 .391
Bloomz-7B1 .461 .500 .579 .367 .396
Bloom-7B1 .431 .501 .573 .262 .387
Teuken-7B-Instruct (Ours) .592 .634 .710 .476 .550
Teuken-7B-Base (Ours) .535 .602 .673 .466 .399
Table 13: Task accuracies for the French language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .616 .635 .649 .614 .565
Mistral-7B-Instruct-v0.3 .590 .625 .627 .545 .565
Occiglot-7B-eu5-Instruct .589 .647 .708 .519 .483
Meta-Llama-3.1-8B .588 .624 .655 .592 .483
Pharia-1-LLM-7B-ctr-aligned .575 .627 .679 .480 .515
Occiglot-7B-eu5 .575 .629 .702 .517 .451
Aya-23-8B .565 .596 .663 .513 .488
Mistral-7B-v0.3 .561 .601 .618 .556 .471
Salamandra-7B-Instruct .558 .626 .658 .472 .478
Pharia-1-LLM-7B-ctr .554 .616 .670 .457 .473
Salamandra-7B .536 .610 .656 .432 .448
Bloomz-7B1 .368 .339 .380 .326 .428
Bloom-7B1 .363 .345 .390 .263 .453
Teuken-7B-Instruct (Ours) .588 .634 .688 .474 .556
Teuken-7B-Base (Ours) .529 .590 .646 .475 .404
Table 14: Task accuracies for the Italian language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .618 .641 .673 .627 .530
Mistral-7B-Instruct-v0.3 .602 .655 .654 .548 .552
Meta-Llama-3.1-8B .597 .632 .679 .606 .470
Occiglot-7B-eu5-Instruct .591 .656 .729 .520 .458
Pharia-1-LLM-7B-ctr-aligned .578 .648 .693 .480 .492
Occiglot-7B-eu5 .576 .636 .712 .516 .438
Mistral-7B-v0.3 .574 .628 .651 .561 .456
Salamandra-7B-Instruct .566 .643 .687 .483 .450
Aya-23-8B .566 .608 .677 .523 .456
Pharia-1-LLM-7B-ctr .552 .636 .681 .453 .437
Salamandra-7B .549 .634 .686 .434 .443
Bloomz-7B1 .478 .517 .582 .379 .434
Bloom-7B1 .437 .509 .573 .260 .407
Teuken-7B-Instruct (Ours) .597 .647 .710 .474 .557
Teuken-7B-Base (Ours) .541 .620 .666 .473 .404
Table 15: Task accuracies for the Spanish language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .608 .625 .651 .622 .535
Mistral-7B-Instruct-v0.3 .590 .625 .625 .551 .558
Meta-Llama-3.1-8B .587 .624 .662 .596 .466
Aya-23-8B .570 .608 .674 .517 .479
Pharia-1-LLM-7B-ctr-aligned .563 .611 .661 .471 .508
Mistral-7B-v0.3 .562 .614 .624 .561 .449
Salamandra-7B-Instruct .557 .631 .667 .473 .459
Pharia-1-LLM-7B-ctr .542 .603 .649 .448 .468
Occiglot-7B-eu5-Instruct .541 .593 .623 .490 .457
Salamandra-7B .535 .619 .669 .430 .422
Occiglot-7B-eu5 .526 .578 .615 .491 .420
Bloomz-7B1 .455 .490 .555 .373 .403
Bloom-7B1 .433 .490 .553 .255 .434
Teuken-7B-Instruct (Ours) .587 .636 .699 .476 .538
Teuken-7B-Base (Ours) .533 .593 .655 .473 .412
Table 16: Task accuracies for the Portuguese language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
Meta-Llama-3.1-8B-Instruct .583 .589 .589 .596 .558
Meta-Llama-3.1-8B .568 .577 .597 .580 .518
Mistral-7B-Instruct-v0.3 .563 .584 .550 .519 .597
Aya-23-8B .559 .578 .638 .505 .516
Mistral-7B-v0.3 .546 .558 .550 .537 .540
Salamandra-7B-Instruct .546 .598 .629 .465 .492
Salamandra-7B .533 .595 .634 .418 .485
Occiglot-7B-eu5-Instruct .490 .516 .499 .433 .510
Occiglot-7B-eu5 .474 .494 .488 .438 .477
Pharia-1-LLM-7B-ctr-aligned .377 .322 .336 .352 .497
Pharia-1-LLM-7B-ctr .372 .314 .333 .350 .490
Bloomz-7B1 .346 .287 .302 .299 .497
Bloom-7B1 .341 .287 .302 .258 .518
Teuken-7B-Instruct (Ours) .578 .592 .650 .470 .601
Teuken-7B-Base (Ours) .525 .558 .607 .457 .479
Table 17: Task accuracies for the Romanian language.
@@
Teuken-7B-Instruct (Ours) .567 .602 .649 .458 .557
Teuken-7B-Base (Ours) .518 .562 .609 .450 .453
@@
Teuken-7B-Instruct (Ours) .578 .597 .692 .471 .555
Teuken-7B-Base (Ours) .528 .566 .642 .467 .437
@@
Teuken-7B-Instruct (Ours) .541 .566 .653 .367 .579
Teuken-7B-Base (Ours) .515 .548 .620 .429 .463
@@
Teuken-7B-Instruct (Ours) .535 .536 .595 .440 .570
Teuken-7B-Base (Ours) .485 .501 .552 .441 .448
@@
Teuken-7B-Instruct (Ours) .550 .559 .631 .447 .564
Teuken-7B-Base (Ours) .493 .520 .590 .446 .415
@@
Teuken-7B-Instruct (Ours) .548 .546 .614 .447 .585
Teuken-7B-Base (Ours) .501 .526 .575 .434 .471
@@
Bloom-7B1 .322 .261 .293 .259 .476
Teuken-7B-Instruct (Ours) .543 .538 .605 .439 .591
Teuken-7B-Base (Ours) .494 .508 .556 .430 .482
@@
Teuken-7B-Instruct (Ours) .537 .520 .585 .425 .619
Teuken-7B-Base (Ours) .495 .492 .549 .427 .514
@@
Teuken-7B-Instruct (Ours) .584 .599 .685 .472 .582
Teuken-7B-Base (Ours) .538 .577 .644 .462 .468
@@
Mistral-7B-Instruct-v0.3 .563 .591 .572 .496 .593
Meta-Llama-3.1-8B-Instruct .547 .554 .544 .546 .546
Salamandra-7B-Instruct .545 .595 .642 .449 .493
Mistral-7B-v0.3 .538 .563 .568 .488 .531
Meta-Llama-3.1-8B .534 .540 .559 .516 .520
Salamandra-7B .519 .586 .638 .389 .463
Occiglot-7B-eu5-Instruct .464 .470 .480 .369 .538
Occiglot-7B-eu5 .448 .464 .474 .344 .511
Aya-23-8B .422 .386 .421 .403 .477
Pharia-1-LLM-7B-ctr .335 .264 .305 .270 .500
Pharia-1-LLM-7B-ctr-aligned .333 .253 .310 .263 .507
Bloom-7B1 .332 .259 .295 .259 .517
Bloomz-7B1 .326 .260 .292 .251 .500
Teuken-7B-Instruct (Ours) .557 .581 .661 .387 .597
Teuken-7B-Base (Ours) .515 .545 .614 .410 .492
Table 27: Task accuracies for the Bulgarian language.
Model Average EU21-ARC EU21-HeSw EU21-MMLU EU21-TQA
@@
Teuken-7B-Instruct (Ours) .576 .589 .656 .456 .603
Teuken-7B-Base (Ours) .518 .555 .606 .444 .468
@@
Teuken-7B-Instruct (Ours) .557 .569 .635 .454 .570
Teuken-7B-Base (Ours) .505 .537 .591 .451 .442
@@
Teuken-7B-Instruct (Ours) .551 .557 .615 .452 .580
Teuken-7B-Base (Ours) .510 .530 .576 .446 .487